Dave was to inherit 5x as much as Libby. Libby was to inherit 2/5 as much as Hector. The sum that would be shared was $102,000. How much will they each get?
Answer by checkley71(8403)

L=2H/5 D=5L OR D=5*2H/5 THUS H+2H/5+5*2H/5=102,000 OR H+2H/5+10H/5=102,000 OR
H+.4H+2H=102,000 OR 3.4H=102,000 OR H=30,000 THEN L=2*30,000/5 OR L=12,000
THEN D=5*12,000=60,000
PROOF 30,000+12,000+60,000=102,000 OR 102,000=102,000